# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v1

@@ Abstract
Accurate temporal extrapolation presents a fundamental challenge for neural operators in modeling dynamical systems, where reliable predictions must extend significantly beyond the training time horizon. Conventional deep operator network (DeepONet) approaches employ two inherently limited training paradigms: fixed-horizon rollouts that predict complete spatiotemporal solutions while disregarding temporal causality, and autoregressive formulations that accumulate errors through sequential predictions. We introduce TI-DeepONet, a framework that integrates neural operators with adaptive numerical time-stepping techniques to preserve the Markovian structure of dynamical systems while substantially mitigating error propagation in extended temporal forecasting. Our approach reformulates the learning objective from direct state prediction to the approximation of instantaneous time-derivative fields, which are subsequently integrated using established numerical schemes. This architecture inherently supports continuous-time prediction and enables the deployment of higher-precision integrators during inference than those utilized during the training phase, balancing computational efficiency with predictive accuracy. We further develop TI(L)-DeepONet, which incorporates learnable coefficients for intermediate slopes in the numerical integration process, thereby adapting to solution-specific variations and enhancing predictive fidelity. Extensive evaluation across three canonical partial differential equations demonstrates that TI(L)-DeepONet marginally outperforms TI-DeepONet, with both methodologies significantly reducing relative $L_{2}$ extrapolation errors, approximately 81% compared to autoregressive implementations and 70% compared to fixed-horizon approaches. Notably, both frameworks maintain prediction stability for temporal domains extending to roughly twice the training interval. This research establishes a physics-aware operator learning paradigm that effectively bridges neural approximation with numerical analysis principles while preserving the causal structure of dynamical systems, addressing a critical capability gap in the long-term forecasting of complex physical phenomena.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) are typically solved using spatial discretization (e.g., finite elements or volumes) combined with time-stepping schemes such as Runge–Kutta or Adams–Bashforth. While accurate, such solvers can be prohibitively expensive for high-dimensional or many-query problems. Explicit integrators are constrained by stability conditions (e.g., CFL), while implicit schemes require solving large systems of equations at each step. Moreover, classical solvers generalize poorly: each new initial or boundary condition requires a fresh simulation.
@@
Neural operators (NOs) offer an attractive alternative by learning mappings from input functions (e.g., initial conditions) to solutions, bypassing discretization and enabling fast inference. Models like DeepONet [1] and FNO [2] have shown promise in parametric PDE settings. However, extending NOs to time-evolving systems introduces new challenges. The standard approach is to train a model that maps the solution at time $t$ to the solution at time $t+\Delta t$, and roll this forward autoregressively (see Figure 1). Unfortunately, this approach suffers from severe error accumulation, as each prediction feeds into the next, compounding inaccuracies and often leading to instability. Alternatively, one can train a model to predict an entire trajectory (“full rollout”), which avoids stepwise feedback but tends to degrade in accuracy when extrapolating beyond the training time window. Crucially, most existing works [3, 2, 4, 1] only evaluate models within the training time span, neglecting the need for reliable extrapolation, a capability that is essential for real-time engineering applications and control tasks. Moreover, existing work [5, 6, 7] on autoregressive and full rollout operator models tends to impose a discrete-time Markovian assumption, which misaligns with the underlying continuous-time dynamics of PDEs. These limitations often result in unstable or inaccurate long-term predictions. Thus, a key barrier remains: NOs struggle not only with long-horizon accuracy but also with extrapolating into unseen temporal domains, a prerequisite for trustworthy deployment.
@@
Several recent works attempt to address these limitations. For example, Diab et al. [8] augment DeepONet with an additional temporal branch to capture dependencies, while Michalowska et al. [9] employ recurrent networks (RNNs, LSTMs, GRUs) atop neural operator outputs to better handle sequences. Other approaches introduce memory modules [10, 11, 12] to compensate for the non-Markovian nature of rollout-based methods. However, these strategies largely rely on generic time-series architectures rather than tackling the issue from a dynamical systems perspective.
@@
In this work, we introduce a novel approach inspired by classical numerical analysis. Our proposed TI-DeepONet framework embeds time integration directly into the operator learning process for dynamical systems. Rather than predicting future solution states directly, we train a DeepONet to approximate the instantaneous time derivative $\partial u/\partial t$. During training, this derivative operator is coupled with a Runge-Kutta numerical integrator to evolve the state forward in time. The predicted future state is then compared against the ground truth in the loss function, enabling the network to learn dynamics that respect temporal causality and numerical stability. At inference time, we leverage high-order multistep integrators (such as Adams-Bashforth/Adams-Moulton predictor-corrector schemes) with refined timesteps to reliably propagate solutions based on the learned time derivative operator. This operator learning framework which maps between function spaces rather than from function spaces to vector spaces as in standard neural networks and neural ODEs [13] - provides the flexibility to modify time-stepping parameters between training and testing phases, thereby ensuring stable predictions. We further introduce TI(L)-DeepONet, an enhanced variant that incorporates learnable weighting coefficients conditioned on the system state at each timestep. This adaptation dynamically adjusts the integration process based on local solution characteristics.
@@
We demonstrate the effectiveness of our approach on three benchmark PDEs – the 1D and 2D Burgers’ equations and the 1D Korteweg–de Vries (KdV) equation. Across all cases, TI-DeepONet and its adaptive variant TI(L)-DeepONet significantly outperform standard autoregressive and full rollout operator models, achieving stable long-term predictions and accurate extrapolation for time horizons approximately twice as long as the training interval.
@@ 2 Methodology
This work uses DeepONet to study error accumulation in autoregressive time-stepping. To mitigate this, we propose TI-DeepONet and TI(L)-DeepONet, which significantly reduce error growth and enable accurate long-term extrapolation. The following subsections introduce these architectures.
@@ 2.1 Deep Operator Network (DeepONet)
Deep Operator Network (DeepONet) are grounded in the universal approximation theorem for nonlinear operators [14], and are designed to learn mappings between infinite-dimensional input and output function spaces. A standard DeepONet architecture comprises two subnetworks: a branch network and a trunk network. The branch network encodes the input function $\mathbf{v}(\eta)$, evaluated at a set of fixed sensor locations $\{\eta_{1},\eta_{2},\dots,\eta_{m}\}$, producing a set of coefficient weights. The trunk network takes as input the query coordinates $\zeta=(x,y,z,t)$ and outputs the corresponding basis functions.
@@ 2.3 Learnable Time Integrator embedded Deep Operator Network (TI(L)-DeepONet)
To introduce flexibility into the integration process, we extend TI-DeepONet by making the Runge–Kutta weighting coefficients learnable and state-dependent. This leads to the development of TI(L)-DeepONet. Specifically, we replace the constant RK4 slope coefficients, $\{\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}\}$, with adaptive coefficients conditioned on the current solution state $u^{i}$. We introduce an auxiliary neural network, $\text{MLP}_{\text{RK}}$, which maps the current state to a set of four weights, $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]=\text{MLP}_{\text{RK}}(u^{i})$. These weights are then used to compute the next solution state via:
@@
For numerical stability and interpretability, we normalize the weights using softmax activation function, such that: $\tilde{\alpha}_{j}=\frac{exp(\alpha_{j})}{\sum_{\ell=1}^{4}exp(\alpha_{\ell})},\forall j=1,2,3,4$. This adaptive scheme allows the model to adjust the influence of each intermediate slope dynamically, thereby compensating for the approximation error [15] of the DeepONet prediction, enabling more robust predictions in stiff, nonlinear, or highly transient regimes. For example, in regions with rapid spatial changes, the model may place greater weight on later-stage evaluations like $k_{4}$, while in smoother regions, earlier evaluations may suffice.
@@ 3 Results
To evaluate the efficacy of our framework, we consider three canonical PDEs: (1) 1D Burgers’ equation, (2) 1D KdV equation, and (3) 2D Burgers’ equation. Model performance is assessed using the relative $L_{2}$ error, defined as $\text{error}=\frac{\|u_{\text{pred}}-u_{\text{true}}\|_{2}}{\|u_{\text{true}}\|_{2}}$, where $u_{\text{pred}}$ and $u_{\text{true}}$ denote the predicted and ground truth solutions, respectively. Table 1 and Figure 3 summarize the comparative training and extrapolation accuracy across all methods for all applications. To evaluate the statistical variability of the performance of the model, we report the mean and standard deviation of error metric in Table A2 based on five independent training trials. Additional details on data generation, training configurations, and network architecture are provided in Supplementary Information (SI) C. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/TI-DeepONet-for-Stable-Long-Term-Extrapolation.git.
Table 1: Relative $L_{2}$ errors across frameworks in the extrapolation regime. TI-DeepONet AB employs RK4 during training and AB2/AM3 during inference.
Problem $t_{train}^{\ast}$ $\Delta t_{e}^{\ast}$ Method Relative $L_{2}$ error
@@
Burgers’ (1D) 0.5 0.01 TI(L)-DeepONet [Ours] 0.0204 0.0243 0.0377 0.0462
TI-DeepONet AB [Ours] 0.0264 0.0310 0.0473 0.0579
Full Rollout 0.0433 0.0965 0.2413 0.3281
Autoregressive 0.5898 0.8742 1.4682 1.7154
KdV (1D) 2.5 0.05 TI(L)-DeepONet [Ours] 0.0884 0.1133 0.1295 0.2010
TI-DeepONet AB [Ours] 0.0861 0.1114 0.1330 0.1941
Full Rollout 0.7769 0.7163 0.7197 0.7951
Autoregressive 0.8127 0.8985 0.9691 1.0626
Burgers’ (2D) 0.33 0.01 TI(L)-DeepONet [Ours] 0.1093 0.1202 0.1419 0.1531
TI-DeepONet AB [Ours] 0.1238 0.1361 0.1609 0.1736
Full Rollout 0.1275 0.1907 0.3649 0.4733
Autoregressive 0.4969 0.5801 0.7604 0.8617
@@ 3.1 One-dimensional Burgers’ Equation
In the full rollout setting, DeepONet learns a mapping $\mathcal{G}_{\boldsymbol{\theta}}\colon s(x)\mapsto u(x,t)$ using a branch network (for the IC) and a trunk network (for spatiotemporal coordinates). A single forward pass predicts the entire solution field. In the autoregressive setting, the models - Autoregressive DeepONet, TI-DeepONet, and TI(L)-DeepONet - the branch network takes as input the state of the system at the current timestep, and the trunk network takes only the spatial discretization. Autoregressive DeepONet predicts the solution at the next time step, while TI-DeepONet and TI(L)-DeepONet predict the time derivative of the solution at the current timestep and integrate temporal dynamics using numerical solvers. TI-DeepONet employs a Heun-type AB2/AM3 predictor–corrector scheme, defined as:
[…]
@@
Figure 4 illustrates the performance of all models for a representative test sample. All models perform well in the training interval; however, the autoregressive DeepONet accumulates error early and diverges during extrapolation. The quantitative trends presented in Table 1 underscore the qualitative observations from Figure 4 as well as Figure 3(a) and (b). The autoregressive DeepONet deteriorates rapidly during extrapolation due to cumulative errors without stabilization mechanisms. While full rollout DeepONet performs better due to its non-recursive architecture, it fails to maintain accuracy beyond the training interval since it does not leverage the Markovian structure of time-dependent PDEs.

### v2

@@ Abstract
Accurate temporal extrapolation presents a fundamental challenge for neural operators in modeling dynamical systems, where reliable predictions must extend significantly beyond the training time horizon. Conventional deep operator network (DeepONet) approaches rely on two inherently limited training paradigms: fixed-horizon rollouts, which predict complete spatiotemporal solutions while disregarding temporal causality, and autoregressive formulations, which accumulate errors through sequential predictions in time. We introduce TI-DeepONet, a framework that integrates neural operators with adaptive numerical time-stepping techniques to preserve the underlying Markovian structure of dynamical systems while substantially mitigating error propagation in extended temporal forecasting. Our approach reformulates the learning objective from direct state prediction to the approximation of instantaneous time-derivative fields, which are subsequently integrated using established numerical schemes. This design naturally supports continuous-time prediction and enables the use of higher-precision integrators at inference than those employed during training, striking a balance between computational efficiency and predictive accuracy. We further develop TI(L)-DeepONet, which incorporates learnable coefficients for intermediate slopes in a multi-stage numerical integration scheme, thereby adapting to solution-specific variations and enhancing predictive fidelity. Rigorous evaluation across four canonical partial differential equations (PDEs) spanning diverse, high-dimensional, chaotic, dissipative, and dispersive dynamics shows that TI(L)-DeepONet marginally outperforms TI-DeepONet, with both frameworks achieving significant reductions in relative $L_{2}$ extrapolation error: approximately 81% compared to autoregressive implementations and 70% compared to fixed-horizon approaches. Notably, both methods maintain stable predictions over temporal domains extending to nearly twice the training interval. This research establishes a physics-aware operator learning paradigm that bridges neural approximation with numerical analysis principles, preserving the causal structure of dynamical systems while addressing a critical gap in long-term forecasting of complex physical phenomena.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) are typically solved using spatial discretization (e.g., finite element or volume) combined with time-stepping schemes such as Runge–Kutta or Adams–Bashforth. While accurate, such solvers can be prohibitively expensive for high-dimensional or many-query problems. Explicit integrators are constrained by stability conditions (e.g., CFL), while implicit schemes require solving large systems of equations at each step. Moreover, classical solvers generalize poorly: each new initial or boundary condition requires a fresh simulation.
@@
Neural operators (NOs) offer an attractive alternative by learning mappings from input functions (e.g., initial conditions) to solutions, bypassing discretization and enabling fast inference. Models like DeepONet [lu2021learning] and FNO [lifourier] have shown promise in parametric PDE settings. However, extending NOs to time-evolving systems introduces new challenges. The standard approach is to train a model that maps the solution at time $t$ to the solution at time $t+\Delta t$, and roll this forward autoregressively (see Figure 1). Unfortunately, this approach suffers from severe error accumulation, as each prediction feeds into the next, compounding inaccuracies and often leading to instability. Alternatively, one can train a model to predict an entire trajectory (“full rollout”), which avoids stepwise feedback but tends to degrade in accuracy when extrapolating beyond the training time window. Crucially, most existing works [lin2023learning, lifourier, wang2023long, lu2021learning] only evaluate models within the training time span, neglecting the need for reliable extrapolation, a capability that is essential for real-time engineering applications and control tasks. Moreover, existing work [xu2023transfer, diab2024temporal, liu2022deeppropnet] on autoregressive and full rollout operator models tends to impose a discrete-time Markovian assumption, which misaligns with the underlying continuous-time dynamics of PDEs. These limitations often result in unstable or inaccurate long-term predictions. Thus, a key barrier remains: NOs struggle not only with long-horizon accuracy but also with extrapolating into unseen temporal domains, a prerequisite for trustworthy deployment.
@@
Several recent works attempt to address these limitations. For example, Diab et al. [diab2025temporal] augment DeepONet with an additional temporal branch to capture dependencies, while Michalowska et al. [michalowska2024neural] employ recurrent networks (RNNs, LSTMs, GRUs) atop neural operator outputs to better handle sequences. Other approaches introduce memory modules [buitragobenefits, he2024sequential, hu5149007deepomamba] to compensate for the non-Markovian nature of rollout-based methods. However, these strategies largely rely on generic time-series architectures rather than tackling the issue from a dynamical systems perspective.
@@
In this work, we introduce a novel approach inspired by classical numerical analysis. Our proposed TI-DeepONet framework embeds time integration directly into the operator learning process for dynamical systems. Rather than predicting future solution states directly, we train a DeepONet to approximate the instantaneous time derivative $\partial u/\partial t$. During training, this derivative operator is coupled with a Runge-Kutta numerical integrator to evolve the state forward in time. The predicted future state is then compared against the ground truth in the loss function, enabling the network to learn dynamics that respect temporal causality and numerical stability. At inference time, we leverage high-order multistep integrators (such as Adams-Bashforth/Adams-Moulton predictor-corrector schemes) with refined timesteps to reliably propagate solutions based on the learned time derivative operator. This operator learning framework which maps between function spaces rather than from function spaces to vector spaces as in standard neural networks and neural ODEs [chen2018neural] - provides the flexibility to modify time-stepping parameters between training and testing phases, thereby ensuring stable predictions. We further introduce TI(L)-DeepONet, an enhanced variant that incorporates learnable weighting coefficients conditioned on the system state at each timestep. This adaptation dynamically adjusts the integration process based on local solution characteristics.
@@
We demonstrate the effectiveness of our approach on four benchmark PDEs: (1) 1D Burgers’, (2) 1D Korteweg-de Vries (KdV), (3) 1D Kuramuto-Sivashinsky (KS), and (4) 2D inviscid Burgers’ equations. We compare our methods against four standard neural operator (NO) baselines: (1) autoregressive DeepONet (DON AR), (2) full rollout DeepONet (DON FR), (3) autoregressive FNO (FNO AR), and (4) full rollout FNO (FNO FR). Across all cases, TI-DeepONet and its adaptive variant TI(L)-DeepONet significantly outperform standard autoregressive and full rollout operator models, achieving stable long-term predictions and accurate extrapolation for time horizons approximately twice as long as the training interval.
@@ 2 Methodology
This study examines the DeepONet framework to investigate error accumulation in autoregressive time-stepping, which serves as the primary motivation for this work. However, it should be noted that this error accumulation phenomenon is not limited to DeepONet but also occurs in other neural operator frameworks, such as FNO. To address this issue, we propose two novel architectures: TI-DeepONet and TI(L)-DeepONet. As will be shown later in this study in Sec. 3, these models effectively tackle this exact issue of autoregressive error accumulation. They are observed to significantly stabilize error growth and enable accurate long-term extrapolation. The following subsections provide a detailed introduction to these architectures.
@@ 2.1 Deep operator network (DeepONet)
Deep operator network (DeepONet) are grounded in the universal approximation theorem for nonlinear operators [chen1995universal], and are designed to learn mappings between infinite-dimensional input and output function spaces. A standard DeepONet architecture comprises two subnetworks: a branch network and a trunk network. The branch network encodes the input function $\mathbf{v}(\eta)$, evaluated at a set of fixed sensor locations $\{\eta_{1},\eta_{2},\dots,\eta_{m}\}$, producing a set of coefficient weights. The trunk network takes as input the query coordinates $\zeta=(x,y,z,t)$ and outputs the corresponding basis functions.
@@ 2.3 Learnable Time Integrator embedded deep operator network (TI(L)-DeepONet)
To introduce flexibility into the integration process, we extend TI-DeepONet by making the Runge–Kutta weighting coefficients learnable and state-dependent. This leads to the development of TI(L)-DeepONet. Specifically, we replace the constant RK4 slope coefficients, $\{\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}\}$, with adaptive coefficients conditioned on the current solution state $u^{i}$. We introduce an auxiliary neural network, $\text{NN}_{\text{RK}}$, which maps the current state to a set of four weights, $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]=\text{NN}_{\text{RK}}(u^{i})$. These weights are then used to compute the next solution state via an adaptive RK4 update:
@@
For numerical stability and interpretability, we normalize the weights using softmax activation function, such that: $\tilde{\alpha}_{j}=\frac{exp(\alpha_{j})}{\sum_{\ell=1}^{4}exp(\alpha_{\ell})},\forall j=1,2,3,4$. This adaptive scheme allows the model to adjust the influence of each intermediate slope dynamically, thereby compensating for the approximation error [lanthaler2022error] of the DeepONet prediction, enabling more robust predictions in stiff, nonlinear, or highly transient regimes. For example, in regions with rapid spatial changes, the model may place greater weight on later-stage evaluations like $k_{4}$, while in smoother regions, earlier evaluations may suffice.
@@ 3 Results
To evaluate the efficacy of our framework, we consider four canonical PDEs: (1) 1D Burgers’ equation, (2) 1D KdV equation, (3) 1D KS equation, and (3) 2D inviscid Burgers’ equation. Model performance is assessed using the relative $L_{2}$ error, defined as $\text{error}=\frac{\|u_{\text{pred}}-u_{\text{true}}\|_{2}}{\|u_{\text{true}}\|_{2}}$, where $u_{\text{pred}}$ and $u_{\text{true}}$ denote the predicted and ground truth solutions, respectively. Table 1 and Figure 3 summarize the comparative training and extrapolation accuracy across all methods for all applications. To evaluate the statistical variability of the performance of the model, we report the mean and standard deviation of error metric in Table A2 based on five independent training trials. Additional details on data generation, training configurations, and network architecture are provided in Supplementary Information (SI) C. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/TI-DeepONet-for-Stable-Long-Term-Extrapolation.git.
Table 1: Relative $L_{2}$ errors across frameworks in the extrapolation regime. TI-DeepONet AB employs RK4 during training and AB2/AM3 during inference. The bold values indicate the best-performing method at each time step, while the underlined values indicate the second-best.
Problem $t_{train}^{\ast}$ $\Delta t_{e}^{\ast}$ Method Relative $L_{2}$ error (in extrapolation)
@@
Burgers’ (1D) 0.5 0.01 TI(L)-DON [Ours] 0.0204 0.0243 0.0377 0.0462
TI-DON AB [Ours] 0.0264 0.0310 0.0473 0.0579
DON Full Rollout 0.0433 0.0965 0.2413 0.3281
DON Autoregressive 0.5898 0.8742 1.4682 1.7154
FNO Full Rollout 0.2298 0.2521 0.2721 0.3551
FNO Autoregressive 0.2377 0.2636 0.2790 0.2771
KdV (1D) 2.5 0.05 TI(L)-DON [Ours] 0.0522 0.0612 0.0701 0.1017
TI-DON AB [Ours] 0.0861 0.1114 0.1330 0.1941
DON Full Rollout 0.7769 0.7163 0.7197 0.7951
DON Autoregressive 0.8127 0.8985 0.9691 1.0626
FNO Full Rollout 0.8666 0.8040 0.7371 0.6352
FNO Autoregressive 0.4171 0.4954 0.5501 0.5931
KS (1D) 15 0.3 TI(L)-DON [Ours] 0.0445 0.079 0.2056 0.3013
TI-DON AB [Ours] 0.0589 0.1066 0.2481 0.3366
DON Full Rollout 0.8298 0.8917 0.8482 0.9073
DON Autoregressive 1.2744 1.2804 1.3204 1.3463
FNO Full Rollout 0.9216 0.9468 0.8943 0.9292
FNO Autoregressive 1.2246 1.2935 1.3207 1.3079
Inviscid Burgers’ (2D) 0.33 0.01 TI(L)-DON [Ours] 0.1838 0.2204 0.2889 0.3207
TI-DON AB [Ours] 0.1911 0.2288 0.3003 0.3341
DON Full Rollout 0.3860 0.4406 0.5019 0.5232
DON Autoregressive 0.7053 0.8589 1.1837 1.3495
FNO Full Rollout 0.3095 0.3472 1.8104 2.2611
FNO Autoregressive 0.0876 0.1729 0.4856 0.7214
@@ 3.1 One-dimensional Burgers’ Equation
In the full rollout setting, DeepONet learns a mapping $\mathcal{G}_{\boldsymbol{\theta}}\colon s(x)\mapsto u(x,t)$ using a branch network (for the IC) and a trunk network (for spatiotemporal coordinates). A similar operator learning problem is defined for FNO. In this case, the mapping is also from the initial condition $s(x)$ to the full spatiotemporal solution field $u(x,t)$. The Fourier blocks in FNO employ an FNO2D-style architecture, applying the Fourier transform across both the $x$ and $t$ coordinates. A single forward pass predicts the entire solution field. In the autoregressive setting, the models - Autoregressive DeepONet, TI-DeepONet, and TI(L)-DeepONet - the branch network takes as input the state of the system at the current timestep, and the trunk network takes only the spatial discretization. Autoregressive DeepONet predicts the solution at the next time step, while TI-DeepONet and TI(L)-DeepONet predict the time derivative of the solution at the current timestep and integrate temporal dynamics using numerical solvers. TI-DeepONet employs a Heun-type AB2/AM3 predictor–corrector scheme, defined as:
$\hat{u}^{n+1}$ $=u^{n}+\Delta t\left[\frac{3}{2}\mathcal{F}(t^{n},u^{n})-\frac{1}{2}\mathcal{F}(t^{n-1},u^{n-1})\right],\quad\text{(Predictor)}$ (4)
$u^{n+1}$ $=u^{n}+\Delta t\left[\frac{5}{12}\mathcal{F}(t^{n+1},\hat{u}^{n+1})+\frac{8}{12}\mathcal{F}(t^{n},u^{n})-\frac{1}{12}\mathcal{F}(t^{n-1},u^{n-1})\right]\quad\text{(Corrector)}$ (5)
Lastly, in accordance with the autoregressive DeepONet learning, autoregressive FNO also operates in an analogous manner, where the mapping is defined as $\mathcal{G}_{\boldsymbol{\theta}}\colon u(x,t)\mapsto u(x,t+1)$. An FNO1D-style architecture is utilized in this case, applying the Fourier transform only in the $x$ dimension.
@@
Figure 4 illustrates the performance of all models for a representative test sample. All models perform well within the training interval; however, the autoregressive DeepONet accumulates error early and diverges during extrapolation. The quantitative trends presented in Table 1 underscore the qualitative observations from Figure 4 as well as Figures 3(a) and (b). The autoregressive DeepONet deteriorates rapidly during extrapolation due to cumulative errors in the absence of stabilization mechanisms. While the full rollout DeepONet performs better owing to its non-recursive architecture, it fails to maintain accuracy beyond the training interval since it does not leverage the Markovian structure of time-dependent PDEs. Regarding the FNO variants, both full rollout and autoregressive FNO perform reasonably well within the training domain. However, in the extrapolation domain, both methods incur relatively higher errors, albeit lower than those of the autoregressive DeepONet. Notably, the autoregressive FNO does not exhibit the steep error growth observed in the autoregressive DeepONet during extrapolation. The full rollout FNO maintains higher errors throughout the entire temporal domain, which can be attributed to its limited capacity to capture the full spectrum of frequencies in the system, as it is trained on solution information from only half of the temporal domain.